OpenStudy (anonymous):

y=(x+3)^2+6 Please advise how to solve

OpenStudy (blurbendy):

which variable are you solving for? y?

OpenStudy (anonymous):

yes

OpenStudy (blurbendy):

you first have to foil (x + 3)(x + 3) what is (x + 3)(x + 3)

OpenStudy (anonymous):

y =(x-3)^2 right

OpenStudy (calculusxy):

Remember the equation of y=mx+b

OpenStudy (calculusxy):

To figure this problem out you can do the vertical line testing and use the given formula above. That will give you the value of y.

OpenStudy (calculusxy):

In UK though, we use y=mx+c

OpenStudy (anonymous):

okay we use it because of the +?

OpenStudy (calculusxy):

I am not sure for the reasoning in the US.

OpenStudy (calculusxy):

However, I will just tell you what the letters mean in the equation of y=mx+b. y=the vertical axis m=slope x=the horizontal axis b=the "y" intercept

OpenStudy (anonymous):

okay I will try to work the problem out. algebra is my weakest subject

OpenStudy (calculusxy):

No problem. :)